# Lookback Lens: Detecting and Mitigating Contextual Hallucinations in Large Language Models Using Only Attention Maps

### v1

@@ 2.2 Experimental Setup
Training the Lookback Lens requires labels for hallucinated and non-hallucinated examples. To obtain these examples, we first prompt LLaMA-2-7B-Chat (Touvron et al., 2023) to greedy decode responses for 1,000 summarization examples from the CNN/DM dataset (See et al., 2017) and 2,655 QA examples from the Natural Questions (Kwiatkowski et al., 2019) following the setup of Liu et al. (2024). More details are shown in Appendix A. Although being prompted to generate correct responses, the decoded responses will contain both hallucinated and non-hallucinated information as the LLaMA model is still not perfect. Then, we employed GPT-4o (OpenAI, 2024) to verify the truthfulness of these responses and provide span-level annotations on hallucinated segments (detailed prompts in Appendix B). Additionally, we performed a pilot study of human annotation on a subset of 70 examples of the summarization task (details in Appendix C), confirming a 97% consistency rate between GPT-4o annotations and human judgments, and validating the reliability of the automated annotations. We show LLaMA-2-7B-Chat’s results on both tasks, as evaluated by GPT-4o, in Table 1. The results show that the generated summaries from LLaMA-2-7B-Chat still exhibit hallucinations about half of the time, highlighting the challenge of summarization tasks.
Dataset Examples Correct
CNN/DM 1000 49.6%
NQ 2655 67.8%
@@
We compare our detection method against several baselines: 1) Text-based entailment classifier: We fine-tune the DeBERTa-v3-base (He et al., 2021) model on the same dataset of CNN/DM and NQ as a natural language entailment (NLI) task. Additionally, we include the results from a state-of-the-art entailment model (Vectara, 2023) trained on a huge amount of annotated NLI data (see details in Appendix E. 2) Hidden states-based classifier: We train classifiers using the same setting as the Lookback Lens but used input features from the hidden states of LLaMA-2-7B-Chat from its 24th, 28th, and 32nd layers instead of the lookback ratio. This baseline resembles a broad range of existing methods in the literature (Azaria and Mitchell, 2023; Simhi et al., 2024).
@@ 3.2 Experimental Setup
To test the Lookback Lens’s effectiveness at transferring across data distributions for the same task (summarization), we use 1,000 examples sampled from the testing set of XSum. Prior studies (Maynez et al., 2020) indicate that traditional evaluation metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019a) correlated poorly with human evaluation on faithfulness and factuality. Recent studies (Chiang and Lee, 2023; Liu et al., 2023) also show a strong correlation between GPT-4 (OpenAI, 2023) evaluation and human evaluation. Thus, we report the averaged accuracy from the binary judgments of GPT-4o, with the prompts in Appendix B. We also conduct a pilot study for human evaluation on GPT-4o’s judgment in Appendix C, finding that 97% of the GPT-4o judgments are consistent with human judgment.
@@
We use the NQ data from the setup of Liu et al. (2024) we describe in Appendix G and evaluate the best span exact match following Kandpal et al. (2023); Mallen et al. (2023).
@@
We consider a multi-turn conversations setup where the model needs to follow previous chat history. We use MT-bench (Zheng et al., 2024), a multi-turn instruction-following benchmark covering 8 categories. We focus exclusively on generating responses for the second turn and use GPT-3.5’s responses as the default for the first turn. We use GPT-4 to score the model’s answers on a scale of 1 to 10 based on various factors, including helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response.
Additionally, since we are particularly interested in mitigating contextual hallucinations, we further exclude math questions and evaluate the remaining 50 general questions. We specifically instruct GPT-4o to focus on whether the responses are faithful to the chat history (see prompt in Appendix B). We refer to this setup as MT-Bench (hallu.).
@@ 3.3 Main Results
We show our results using 8 candidates per chunk in a chunk size of 8 in Table 3, and the ablation with different chunk sizes is shown in Table 6. Lookback Lens Guided Decoding can improve the performance on both in-domain task (XSum, by 9.6%) and out-of-domain tasks (NQ, by 3%). The original greedy decoding results on XSum achieved 49.0% correct which means 510 examples were hallucinated. Our decoding method significantly reduced the number of hallucinated examples from 510 to 414, resulting in an 18.8% reduction in the hallucinated examples. This result is on par with using SoTA NLI to guide the decoding, where SoTA NLI is trained on roughly 731k annotated summarization examples, which is 700$\times$ larger compared to our 1k training set. (See Appendix E.) In contrast, decoding guided by hidden states-based or the NLI (our implementation) classifiers, both trained on the same data of our method, can only slightly improve the performance on NQ, but not for XSum, probably due to the issue of distribution shift, highlighting the advantages of Lookback Lens in generalization ability.
@@ 4 Cross-model Transfer
Since the total numbers of attention heads are different in 7B and 13B models, and there is no obvious one-to-one mapping between the heads, we use a linear regression model to map the heads from the 13B model to the heads in 7B model. Concretely, we have 1024 heads in 7B and 1600 heads in 13B. We extract the averaged lookback ratio per head for all the $|D|$ training examples, resulting in a $1024\times|D|$ matrix and a $1600\times|D|$ matrix. [^2] To ensure that two models are generating the same content when extracting lookback ratio, we decode from 7B and run the 13B model on the 7B outputs. We then fit a linear regression model to map the heads to reconstruct the 7B heads from 13B heads. After applying the linear transformation to the lookback ratio from 13B, the transformed heads can be directly used by 7B’s classifiers. See details in Appendix E.
@@ 5 Discussions and Ablations
More interestingly, we also include the results that only select the top-$k$ heads among the heads with most positive/negative coefficients, which are positive/negatively correlated to factuality. On the heads with positive coefficients, higher lookback ratio (i.e., when the heads attend at the context more) indicates higher factuality and less hallucination; conversely, heads with negative coefficients suggest a lower lookback ratio (i.e., attending to generated tokens more) is more likely to be truthful. Table 7 shows that none of positive or negative heads alone can be on par with using the top-$k$ largest magnitude heads. This result implies that both positive and negative heads are critical for a model to generate factual responses. We conjecture that the positive heads may specialize at context grounding, and thus higher lookback ratio on these heads leads to more factual response. On the other hand, the negative heads may be critical at ensuring consistency in its own generation, and thus should attend to the generated tokens more. We leave further investigation on this interesting balance for future work. Meanwhile, we visualize the lookback ratio of positive/negative heads in Appendix F.
@@ 6 Related Work
Classifier guided generation.
@@
Self-attention and model behavior.
@@ Appendix A Data Creation for Lookback Lens
Appendix B Evaluation Prompt for GPT-4o
@@
This prompt is used for 1) collecting the data to train the Lookback Lens in Table 1, and 2) evaluating the XSum summarization task in Sections 3, 4, and 5.
@@
Appendix C Pilot Study of Human Evaluation on GPT-4o Evaluation
To assess the quality of GPT-4o’s evaluation in Appendix B, we conducted a pilot study using 70 examples from XSum data, involving the authors and colleagues who participated voluntarily as evaluators. Evaluators are all native English speakers. The human evaluators are provided with the document, ground truth summary, summary from LLaMA-2-7B-Chat, and the judgment from GPT-4o. And the task is to give a binary judgment on whether the judgment from GPT-4o is correct. The screenshot of our interface is shown in Figure 4. Every evaluator is informed and consents that the annotation result will only be used to report statistics in this paper. Our pilot study shows that 68 out of 70 (97%) GPT-4o outputs are correct, justifying the use of GPT-4o as an automatic evaluator.
@@ Appendix C Pilot Study of Human Evaluation on GPT-4o Evaluation
Appendix D Evaluation Prompt for MT-Bench
@@
Appendix E Model Details
@@
We run all the models on NVIDIA A6000 (48GB) and V100 (32GB) GPUs. We do not train the model, but only run the inference part. Each of the examples takes around 20-30 seconds for 7B model, 40-60 seconds for 13B model to generate responses using our Lookback Lens Guided Decoding. Please check Appendix G to estimate the total running time on each of the datasets, as it depends on number of examples.
@@
Appendix F Visualization
@@
Appendix G Dataset Details
The datasets we used in the paper have the following details:

### v2

@@ 2.2 Experimental Setup
Training the Lookback Lens requires labels for hallucinated and non-hallucinated examples. To obtain these examples, we first prompt LLaMA-2-7B-Chat (Touvron et al., 2023) to greedy decode responses for 1,000 summarization examples from the CNN/DM dataset (See et al., 2017) and 2,655 QA examples from the Natural Questions (Kwiatkowski et al., 2019) following the setup of Liu et al. (2024). More details are shown in Appendix A. Although being prompted to generate correct responses, the decoded responses will contain both hallucinated and non-hallucinated information as the LLaMA model is still not perfect. Then, we employed GPT-4o (OpenAI, 2024) to verify the truthfulness of these responses and provide span-level annotations on hallucinated segments (detailed prompts in Appendix B.1).
Additionally, we performed a pilot study of human annotation on a subset of 70 examples of the summarization task (details in Appendix B.2), confirming a 97% consistency rate between GPT-4o annotations and human judgments, and validating the reliability of the automated annotations. We show LLaMA-2-7B-Chat’s results on both tasks, as evaluated by GPT-4o, in Table 1. The results show that the generated summaries from LLaMA-2-7B-Chat still exhibit hallucinations about half of the time, highlighting the challenge of summarization tasks.
@@
We compare our detection method against several baselines: 1) Text-based entailment classifier: We fine-tune the DeBERTa-v3-base (He et al., 2021) model on the same dataset of CNN/DM and NQ as a natural language entailment (NLI) task. Additionally, we include the results from a state-of-the-art entailment model (Vectara, 2023) trained on a huge amount of annotated NLI data (see details in Appendix C.1).
2) Hidden states-based classifier: We train classifiers using the same setting as the Lookback Lens but used input features from the hidden states of LLaMA-2-7B-Chat from its 24th, 28th, and 32nd layers instead of the lookback ratio. This baseline resembles a broad range of existing methods in the literature (Azaria and Mitchell, 2023; Simhi et al., 2024). Our selection of layers followed the findings outlined in Azaria and Mitchell (2023), which used layers 32, 28, 24, and 20 of a 32-layer LLM for detecting hallucinations. They find that layers near the 28th layer are most effective (see Table 3 and 4 in Azaria and Mitchell (2023)).
We include additional experiments for leveraging multiple layers or all layers in predicting contextual hallucinations in Appendix D.2, but the results are not significantly better than using the 28th layer. Some papers suggest attention block outputs could be more useful for detecting hallucinations (Campbell et al., 2023; Li et al., 2024), we include the additional comparative experiments in Appendix D.3, but the difference between hidden states and attention block outputs is relatively small.
@@ 3.2 Experimental Setup
To test the Lookback Lens’s effectiveness at transferring across data distributions for the same task (summarization), we use 1,000 examples sampled from the testing set of XSum. Prior studies (Maynez et al., 2020) indicate that traditional evaluation metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019a) correlated poorly with human evaluation on faithfulness and factuality. Recent studies (Chiang and Lee, 2023; Liu et al., 2023) also show a strong correlation between GPT-4 (OpenAI, 2023) evaluation and human evaluation. Thus, we report the averaged accuracy from the binary judgments of GPT-4o, with the prompts in Appendix B.1. We also conduct a pilot study for human evaluation on GPT-4o’s judgment in Appendix B.2, finding that 97% of the GPT-4o judgments are consistent with human judgment.
@@
We use the NQ data from the setup of Liu et al. (2024) we describe in Appendix C.2 and evaluate the best span exact match following Kandpal et al. (2023); Mallen et al. (2023).
@@
We consider a multi-turn conversations setup where the model needs to follow previous chat history. We use MT-bench (Zheng et al., 2024), a multi-turn instruction-following benchmark covering eight categories. We focus exclusively on generating responses for the second turn and use GPT-3.5’s responses as the default for the first turn. We use GPT-4 to score the model’s answers on a scale of 1 to 10 based on various factors, including helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response.
Additionally, since we are particularly interested in mitigating contextual hallucinations, we further exclude math questions and evaluate the remaining 50 general questions. We specifically instruct GPT-4o to focus on whether the responses are faithful to the chat history (see prompt in Appendix B.1). We refer to this setup as MT-Bench (hallu.).
@@ 3.3 Main Results
We show our results using eight candidates per chunk in a chunk size of eight in Table 3, and the ablation with different chunk sizes is shown in Table 6. Lookback Lens Guided Decoding can improve the performance on both in-domain task (XSum, by 9.6%) and out-of-domain tasks (NQ, by 3%). The original greedy decoding results on XSum achieved 49.0% correct which means 510 examples were hallucinated. Our decoding method significantly reduced the number of hallucinated examples from 510 to 414, resulting in an 18.8% reduction in the hallucinated examples. This result is on par with using SoTA NLI to guide the decoding, where SoTA NLI is trained on roughly 731k annotated summarization examples, which is 700$\times$ larger compared to our 1k training set. (See Appendix C.1.) In contrast, decoding guided by hidden states-based or the NLI (our implementation) classifiers, both trained on the same data of our method, can only slightly improve the performance on NQ, but not for XSum, probably due to the issue of distribution shift, highlighting the advantages of Lookback Lens in generalization ability.
@@ 4 Cross-model Transfer
Since the total numbers of attention heads are different in 7B and 13B models, and there is no obvious one-to-one mapping between the heads, we use a linear regression model to map the heads from the 13B model to the heads in 7B model. Concretely, we have 1024 heads in 7B and 1600 heads in 13B. We extract the averaged lookback ratio per head for all the $|D|$ training examples, resulting in a $1024\times|D|$ matrix and a $1600\times|D|$ matrix. [^2] To ensure that two models are generating the same content when extracting lookback ratio, we decode from 7B and run the 13B model on the 7B outputs. We then fit a linear regression model to map the heads to reconstruct the 7B heads from 13B heads. After applying the linear transformation to the lookback ratio from 13B, the transformed heads can be directly used by 7B’s classifiers. See details in Appendix C.1.
@@ 5 Discussions and Ablations
More interestingly, we also include the results that only select the top-$k$ heads among the heads with most positive/negative coefficients, which are positive/negatively correlated to factuality. On the heads with positive coefficients, higher lookback ratio (i.e., when the heads attend at the context more) indicates higher factuality and less hallucination; conversely, heads with negative coefficients suggest a lower lookback ratio (i.e., attending to generated tokens more) is more likely to be truthful. Table 7 shows that none of positive or negative heads alone can be on par with using the top-$k$ largest magnitude heads. This result implies that both positive and negative heads are critical for a model to generate factual responses. We conjecture that the positive heads may specialize at context grounding, and thus higher lookback ratio on these heads leads to more factual response. On the other hand, the negative heads may be critical at ensuring consistency in its own generation, and thus should attend to the generated tokens more. We leave further investigation on this interesting balance for future work. Meanwhile, we visualize the lookback ratio of positive/negative heads in Appendix D.1.
@@ 6 Related Work
Classifier Guided Generation
@@
Self-attention and Model Behavior
@@ 7 Conclusion
Acknowledgement
We sincerely thank Philip Schroeder, Huirong Wen, Andrew Rouditchenko, Nishad Gothoskar, Ani Nrusimha, Howard Chen, Weijia Shi, and Nour Jedidi for their discussion and help in this project. This research was sponsored by the United States Air Force Research Laboratory and the United States Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. Linlu and Yoon were supported in part by MIT-IBM Watson AI Lab.
@@ Appendix A Data Creation for Lookback Lens
Appendix B Evaluation Details
B.1 Evaluation Prompt for GPT-4o
@@
This prompt is used for 1) collecting the data to train the Lookback Lens in Table 1, and 2) evaluating the XSum summarization task in Sections 3, 4, and 5. We also provide the approximate cost of GPT-4o calls (in USD):
  • 1000 examples from XSum is around \$8.
  • 1000 examples from CNN/DM is around \$12.
  • 2655 examples from NQ is around \$16.
@@
B.2 Human Evaluation on GPT-4o Evaluation
Summarization
To assess the quality of GPT-4o’s evaluations, we initially conducted a pilot study using 70 XSum dataset examples, with native English-speaking authors and colleagues as evaluators. Evaluators received the document, ground truth summary, LLaMA-2-7B-Chat’s summary, and GPT-4o’s judgment to provide a binary judgment on GPT-4o’s accuracy. Our interface is depicted in Appendix B.1 (see Figure 4). This initial evaluation affirmed the correctness of GPT-4o’s judgments in 68 out of 70 cases. To further verify these results, we expanded our evaluation through Amazon MTurk, adding two additional annotations per example. Across all 210 evaluations (70 initial + 140 MTurk), only 9 annotations were marked incorrect, and in only 2 cases did a majority of annotators deem the judgment incorrect (marked incorrect by at least two annotators). With a final accuracy of 97.1%, and high intra-annotator agreement, the comprehensive evaluation supports GPT-4o’s use as an automatic evaluator for the entire dataset.
Question Answering
We expand the human evaluation to Natural Questions dataset using Amazon MTurk. The evaluation interface is copied from the summarization setup, but changing “summary” to “answer”, as well as adding the “question” field.
We take 50 examples and assign each example to three different annotators. There are 7 annotations marked incorrect out of the 150 annotations. In total, 3 of the examples are marked incorrect by at least two annotators. If applying a majority vote, 47 out of 50 examples are correct, resulting in a 94.0% accuracy. This suggests that it is generally sufficient to use GPT-4o to verify the generated responses on the question-answering task.
@@ B.2 Human Evaluation on GPT-4o Evaluation
B.3 Evaluation Prompt for MT-Bench
@@
Appendix C Experiment Details
C.1 Model Details
@@
We run all the models on NVIDIA A6000 (48GB) and V100 (32GB) GPUs. We do not train the model, but only run the inference part. Each of the examples takes around 20-30 seconds for 7B model, 40-60 seconds for 13B model to generate responses using our Lookback Lens Guided Decoding. Please check Appendix C.2 to estimate the total running time on each of the datasets, as it depends on number of examples.
@@
C.2 Dataset Details
The datasets we used in the paper have the following details:
Appendix D Additional Results
D.1 Visualization
@@
D.2 Using Multiple or All Layers for Hidden States
Multiple Layer
We follow the prior study (Azaria and Mitchell, 2023) to use the layers with the best predictive power in hallucination detection: 32nd/28th/24th/20th layers. We concatenate the 4 layer features into a huge feature. Please note that the hidden dimension of LLaMA-7B is 4096, so combining 4 layers would result in a 16384-dim feature vector. In contrast, our Lookback Lens feature for the 7B model is only 1024-dim. Thus, the big classifier using 16384 input features is supposed to be more effective given that it uses 10x more features.
However, the result shown in Table 13 indicates that concatenating 4 layers is still less effective compared to our Lookback Lens.
All Layers
We also try to use the hidden states from all layers, but concatenating them all will result in a huge feature vector with dimensions of more than 100k and make the classifier extremely slow in training. Thus, we perform max/average pooling for the features across different layers, resulting in 4096-dim feature vectors as the classifier inputs. The results shown in the table below are still worse than our Lookback Lens results.
The two experiments above indicate that using multiple or all layers may not be the key to making the classifier accurate. Instead, by designing good features like lookback ratio, the compact 1024-dim feature can be even more effective compared to the 10x bigger high-dimensional hidden state features.
Method AUROC (sliding window = 8)
NQ $\rightarrow$ Sum. Sum. $\rightarrow$ NQ
Residual outputs (hidden states)
Layer 32 56.1 59.4
Layer 28 57.7 58.8
Layer 24 58.3 58.3
Layer 20 57.6 59.5
Concatenate above 4 layers 58.8 59.2
Max pooling all 32 layers 56.7 59.2
Average pooling all 32 layers 57.3 59.2
Ours: Lookback Lens 66.1 66.0
Table 13: AUROC results for different methods of utilizing hidden states.
D.3 Comparing Attention Outputs with Hidden States
Some papers mention that attention block outputs could be more useful for detecting hallucinations (Campbell et al., 2023; Li et al., 2024), while our main experiments only consider the hidden states as input features for detecting contextual hallucinations. Here we include additional experiment results that use attention block outputs instead. In Table 14, we show that there is no significant difference when switching to attention block outputs, and our Lookback Lens still outperforms these baselines.
Method AUROC (sliding window = 8)
NQ $\rightarrow$ Sum. Sum. $\rightarrow$ NQ
Attention block outputs
Layer 32 57.6 60.7
Layer 28 58.5 57.2
Layer 24 56.3 57.2
Residual outputs (hidden states)
Layer 32 56.1 59.4
Layer 28 57.7 58.8
Layer 24 58.3 58.3
Ours: Lookback Lens 66.1 66.0
Table 14: AUROC results for different layers and outputs.